OpenStudy (anonymous):

Any idea how to evaluate pi/2 int_{0}^{infty}i ^{(ix)} dx ?

OpenStudy (anonymous):

The integral I'm looking for is: \[\pi/2 \int\limits_{0}^{\infty}i ^{(ix)} dx\]

OpenStudy (anonymous):

just notice \(\huge i=e^{\frac{\pi}{2}i} \)

OpenStudy (anonymous):

Thanks mukushla, I then get \[i^{(ix)} = e^{(\pi/2)x i^2} = e^{-(\pi/2)x} \]

OpenStudy (anonymous):

yep; exactly

OpenStudy (anonymous):

This integral looks a lot easier to solve :)